the measure of the angles of a triangle are 2x + 10 , 3x , and 8x - 25,
the triangle is
A obtuse B right C acute D equilateral E isosceles

Found 2 solutions by CharlesG2, nyc_function:
Answer by CharlesG2(834)

the measure of the angles of a triangle are 2x + 10 , 3x , and 8x - 25,
the triangle is
A obtuse B right C acute D equilateral E isosceles
in a triangle the sums of the angles must be 180
2x + 10 + 3x + 8x - 25 = 180
(added angles together and set equal to 180, now need to solve for x)
13x - 15 = 180
13x = 195
x = 15
angles:
2x + 10 --> 40
3x --> 45
8x - 25 --> 95
no angles equal to 90, no angles equal, largest angle greater than 90
the triangle is A obtuse

Answer by nyc_function(2741)

3x + 2x + 10 + 8x - 25 = 180
13x - 15 = 180
13x = 180 + 15
13x = 195
x = 195/13
x = 15
The three angles are: 3(15) = 45, 2(15) + 10 = 40 and 8(15) - 25 = 95
The BIG angle of 95 degrees points me to answer choice A for obtuse.